OpenStudy (anonymous):

Which of the follow functions are linear? y - 4 = -8(x – 1) y = x4 y – x2 = 4.5 3x + 5y = 15

OpenStudy (anonymous):

@jhonyy9

OpenStudy (anonymous):

@Mehek14

jhonyy9 (jhonyy9):

so do you know what mean ,,function linear" ?

OpenStudy (anonymous):

if it makes a straight line

OpenStudy (wolf1728):

Functions that have exponents (x^2, x^3, etc) are NOT linear. So, y - 4 = -8(x – 1) and 3x + 5y = 15 are linear

OpenStudy (anonymous):

what do u think